OpenStudy (anonymous):

find the integers who reciprocals add up to 9/20 @Abdulhameed

OpenStudy (anonymous):

sorry i haven't understood the question i don't know what reciprocal mean i'm arabic i don't know all the mathematical vocabularies so sorry

OpenStudy (anonymous):

reciprocal means 1/9 will be 9/1 you just switch the fractions around

OpenStudy (anonymous):

aha thanks for the information

OpenStudy (anonymous):

let the first integer is x and the second integer is x+1

OpenStudy (anonymous):

then \[\frac{ 1 }{ x }+\frac{ 1 }{ x+1 }=\frac{ 9 }{ 20 }\]

OpenStudy (anonymous):

5 + 4 = 9 reciprocals : 1/5 + 1/4 = 4/20 + 5/20 = 9/20

OpenStudy (anonymous):

understand ?

OpenStudy (anonymous):

that's true but u have to show how to find them

OpenStudy (anonymous):

oh.....I did the reciprocals first....then did the numbers. I had to find 2 numbers who's common denominator was 20. I guess I did it backwards

OpenStudy (anonymous):

lol me 2 but i was thinking of a way to make it a quadratic equation

OpenStudy (anonymous):

oh...good luck with that. I wouldn't make it more complicated then it is